Solve the problem. If necessary, round to the nearest tenth.
On a sunny day, a tree and its shadow form the sides of a right triangle. If the hypotenuse is 35 m long and the tree is 28 m tall, how long is the shadow?

Answer by checkley77(12844)

x^2+28^2=35^2
x^2+784=1,225
x^2=1,225-784
x^2=441
x=sqrt441
x=21 ans. for the length of the shadow.